OpenStudy (anonymous):

the oscillation of tides in Laguna Beach, California can be modeled with a sinusoidal function. The difference between the highest and lowest point is 7.4 and the period is 12.8 hours. What function represents the position D of the tides of time t.

OpenStudy (anonymous):

the amplitude should be half of \(7.4\) which is \(3.7\)

OpenStudy (anonymous):

it doesn't say whether to use sine or cosine, so lets use sine

OpenStudy (anonymous):

all of the options of cosine

OpenStudy (anonymous):

the period is given at \(12.8\) ok so use \(3.7\cos(bt)\)

OpenStudy (anonymous):

to find \(b\) the period is \(\frac{2\pi}{b}\) so set \[\frac{2\pi}{b}=12.8\]and solve for \(b\)

OpenStudy (anonymous):

you get \(b=\frac{2\pi}{12.8}\) or \(b=\frac{\pi}{6.4}\) unless you want a decimal

OpenStudy (anonymous):

i don't know what your choices are but it should look something like \[3.7\cos(\frac{\pi t}{6.4})\]

OpenStudy (anonymous):

yeah thats the answer. thanks a lot...:D

OpenStudy (anonymous):

yw